### v1

# Beyond Reference-Based Metrics: Analyzing Behaviors of Open LLMs on Data-to-Text Generation

Abstract
We investigate to which extent open large language models (LLMs) can generate coherent and relevant text from structured data. To prevent bias from benchmarks leaked into LLM training data, we collect Quintd-1: an ad-hoc benchmark for five data-to-text (D2T) generation tasks, consisting of structured data records in standard formats gathered from public APIs. We leverage reference-free evaluation metrics and LLMs’ in-context learning capabilities, allowing us to test the models with no human-written references. Our evaluation focuses on annotating semantic accuracy errors on token-level, combining human annotators and a metric based on GPT-4. Our systematic examination of the models’ behavior across domains and tasks suggests that state-of-the-art open LLMs with 7B parameters can generate fluent and coherent text from various standard data formats in zero-shot settings. However, we also show that semantic accuracy of the outputs remains a major issue: on our benchmark, 80% of outputs of open LLMs contain a semantic error according to human annotators (91% according to GPT-4). Our code, data, and model outputs are available at https://d2t-llm.github.io.
1 Introduction
Figure 1: We experiment with using LLMs for generating text from structured data in various domains, analyzing model behavior and evaluating their output on token level.
Task Id Domain Task Description Source Format
openweather Weather Generating a weather forecast from weather data. OpenWeather JSON
gsmarena Technology Describing a product based on its attributes. GSM Arena JSON
ice_hockey Sport Describing an outcome of an ice-hockey game. Rapid API JSON
owid Health Generating a caption for a time series. Our World In Data CSV
wikidata World facts Desribing entities and relations in a knowledge graph. Wikidata Markdown
Table 1: The domains and tasks included in the Quintd data collection tool we use for testing D2T generation with LLM. In our experiments, we download 100 development and 100 test examples of input data for each domain.
Large language models (LLMs; Ouyang et al., 2022; Touvron et al., 2023a, b; Jiang et al., 2023; Tunstall et al., 2023) have already left a mark in many areas of natural language processing (NLP). Suprisingly, their applicability to the task of data-to-text (D2T) generation Reiter and Dale (1997); Gatt and Krahmer (2018) remains underexplored, with limited evaluation on a handful of well-established benchmarks only Axelsson and Skantze (2023); Yuan and Färber (2023). Generating text from structured data is arguably challenging for LLMs, given the specifics of D2T generation such as long inputs, complex non-linear structure, and strict requirements on semantic accuracy. However, a bigger issue is perhaps the lack of testing grounds. The current D2T generation benchmarks are saturated by approaches optimizing towards traditional reference-based evaluation metrics, which were shown to correlate poorly with human judgement Gehrmann et al. (2023); van der Lee et al. (2021); Novikova et al. (2017). When it comes to the models, using closed LLMs OpenAI (2023a, b) is increasingly considered a bad research practice due to its non-reproducibility Rogers (2023); Chen et al. (2023), and potential data contamination further restricts the space for experiments Golchin and Surdeanu (2023); Aiyappa et al. (2023); Anonymous (2023).
In this paper, we propose how to get around these issues and move towards better D2T generation approaches with more rigorous evalution. We begin with the realization that unlabeled data are plentiful. To leverage the data for our experiments, we introduce Quintd: [^1] Quintet of Unlabeled Inputs for Natural Tasks in Data-to-text (pronounced as “quintet”). a tool for collecting structured data from five domains in standard formats: JSON, CSV, and Markdown. We choose the domains so that the data can be directly used as input for five distinct D2T generation tasks. Specifically, our tasks inlude generating weather forecasts, sports reports, product descriptions, chart captions, and entity descriptions (see Table 1). Next, we collect a set of 1000 inputs with Quintd and use the inputs as an ad-hoc benchmark (called Quintd-1) for testing abilities of LLMs for D2T generation. We assume that the data formats in Quintd-1 are common in the LLMs’ pretraining corpora, so we specify the task using instructions (instead of standard finetuning with human-written outputs), capitalizing on the in-context learning abilities of LLMs (§2).
We push towards better reproducibility by focusing on open LLMs, which – apart from being more accessible – also achieve increasingly better results across tasks Zheng et al. (2023); Beeching et al. (2023). For our experiments, we use three open LLMs with 7B parameters: Llama-2 Touvron et al. (2023b); TogetherAI (2023), Mistral Jiang et al. (2023), and Zephyr Tunstall et al. (2023). We also use GPT-3.5 OpenAI (2023b) as a closed model baseline for the final experiments. Given the behavioral nature of the experiments with LLMs Holtzman et al. (2023), we put emphasis on reporting model behavior throughout the process (§3).
Another piece of the puzzle is reference-free evaluation: using the input data as a ground for comparison, instead of human references (§4). For evaluation, we use manual annotations from human crowdworkers van der Lee et al. (2021), along with a customized automatic metric based on GPT-4 Liu et al. (2023); Chiang and Lee (2023); Kocmi and Federmann (2023a). To get a fine-grained picture of model errors, we annotate semantic accuracy errors on the level of individual tokens Thomson and Reiter (2020); Thomson et al. (2023).
Based on our results, we provide general recommendations about D2T generation with open LLMs across tasks and formats (§5). Our main findings are as follows:
  • Open LLMs can generate fluent outputs from structured data in common formats under zero-shot settings.
  • Semantic accuracy is still a major obstacle: on average, 80% of LLM outputs contain a semantic error according to human annotators and 91% according to the GPT-4-based metric.
  • Long data inputs cause practical issues, including the need for long-context models, increased GPU memory requirements, and unavailability of few-shot approaches.
  • The output quality can be aided by following several rules-of-thumb on preprocessing the model input, such as including units, removing unnecessary fields, or prefixing the model answer.
2 Reference-Free D2T Generation
2.1 Data Collection Tool
We introduce a tool dubbed Quintd for collecting ad-hoc tests sets using public APIs in five different domains, and we collect one such set (Quintd-1) for our experiments. Our main reasons for departing from the traditional scheme of benchmarking on well-established datasets are:
  1. 1.
    Any published test sets may be potentially included in the training data of LLMs.
  2. 2.
    Public sources of structured data offer enough resources for creating ad-hoc test sets.
  3. 3.
    Without human references, our data collection scheme is lightweight and replicable.
Given the available public sources of data, we settled on the five tasks which are described in Table 1 (see Appendix A for more details). The tasks are based on structured data in common formats: JSON, CSV, and Markdown.
Prompt
Start with
Figure 2: The prompt $\mathcal{P}$ and the model output prefix (“Start with”) we used for the experiments in this paper. DATA is filled with the data record $x$ and output_type is filled accordingly for each domain $\mathcal{D}$ (cf. Table 1).
2.2 Quintd-1 Dataset
The benchmark we collected using Quintd for our experiments in this paper (Quintd-1) contains 500 examples in the development set and 500 examples in the test set (100 examples per domain for each split). We keep the size of the dataset moderate for a quick experimental turnaround.
We downloaded the data between November 2023 and January 2024. Note that the dataset contains only unlabeled data without any reference outputs (e.g. weather data, but not a textual weather forecast). New versions of the benchmark can be easily created with the Quintd tool we provide.
2.3 Task Definition
Each example in Quintd-1 consists of a structured data record $x$ from a domain $\mathcal{D}\in$openweather, gsmarena, ice_hockey, owid, wikidata}. Given $x$ and a prompt $\mathcal{P}$, the goal is to generate natural language output $y$ faithful to the data $x$, according to the instructions in the prompt $\mathcal{P}$ (see Figure 2).
3 Experiments
3.1 Experimental Process
Our goal is to avoid extensive data preprocessing and prompt engineering, since these steps could harm the reproducibility and generalizability of our experiments. With this goal in mind, we decided to use the same prompt template $\mathcal{P}$ for all the domains and models.
For a set of preliminary experiments, we first wrote down the initial version of the prompt and used the data without further preprocessing. We then iteratively improved our experimental setup by observing outputs on the development set. We describe all the observations and modifications we made before generating the final outputs on the test set in §3.3.
Figure 3: Our experimental setup. We first generate the outputs using LLMs which are given raw data and a task-specific prompt. We annotate the token-level semantic errors in the LLM outputs with a) an automatic metric based GPT-4 which matches the output to the raw data, and b) human annotators, who annotate the errors in the output given the data visualization.
3.2 Models
For our experiments, we selected the following LLMs available under an open license:
  • Llama2 Touvron et al. (2023a); TogetherAI (2023),
    togethercomputer/Llama-2-7B-32K-Instruct
  • Mistral Jiang et al. (2023),
    mistralai/Mistral-7B-Instruct-v0.1
  • Zephyr Tunstall et al. (2023).
    HuggingFaceH4/zephyr-7b-beta
The models are instruction-tuned, operate with 32k context, and have good performance on recent benchmarks. All the models have 7B parameters and thus fit on a single NVIDIA A40 (48G VRAM) in 16-bit precision. The models are available through HuggingFace Wolf et al. (2020).
We accessed the models via an API provided by the text-generation-webui framework [^2] https://github.com/oobabooga/text-generation-webui running locally. For the final experiments, we also included GPT-3.5 OpenAI (2023b) accessed through the OpenAI API.
3.3 Observations from Preliminary Experiments
During development, we made several observations which we took into account for our final experimental setup:
Any input field may appear in the output.
The models do not always select the most relevant fields for the given output. For example, we observed that the models commonly mention identifiers, timestamps, files, and other metadata, leading to unnatural outputs. Due to this, we decided not to include these irrelevant fields on the input.
Units need to be specified explicitly.
If the units are not specified in the data record, the models tend to resort to their best guess. This may go unnoticed if the unit is clear from the context (e.g., the model will usually not report temperature in Fahrenheit instead of Celsius), but it may get problematic if the value is ambiguous (e.g., wind speed in km/h versus m/s). Therefore, we explicitly add units to all data records where appropriate.
Understandable labels are enough.
On the flip side, we decide not to add extra descriptions to the keys if the key was understandable from the label (e.g. homeTeam or dimensions). As discussed by Kasner et al. (2023b), pretrained models tend to interpret the fields correctly as long as the label is human-readable. We only decided to include chart metadata for the CSV files in the owid domain.
Long inputs can be troublesome.
The inputs in some domains can easily get longer than 10-20k tokens. This issue is amplified by the fact that the evaluated LLMs tokenize numbers into individual digits. Because of the long inputs, we decided to give up the official Llama-2-7B-Instruct model with 4k context Touvron et al. (2023b) in favor of Llama2 with 32k context TogetherAI (2023). With long inputs, the GPU memory consumption also gets considerably higher, so we needed to downsample the data in owid and openweather to keep their length under ~8k tokens.
Few-shot experiments are infeasible.
Due to the context-length limitations described above, we were not able to conduct few-shot experiments since we could not robustly fit an additional ($x_{\text{example}}$, $y_{\text{example}}$) pair in the prompt. We attempted including only $y_{\text{example}}$ (making the setup “half-shot”), but we observed that the models then tended to use entities from the example (unrelated to the actual input) in their outputs. Therefore, we decided not to to follow this line of experiments (see §5.3 for discussion).
Deterministic decoding and sampling are on par.
In our preliminary experiments, we observed a roughly similar output quality for both deterministic decoding and sampling. [^3] We used the text-generation-webui default decoding parameters: temperature=0.7, top_p=0.9, and top_k=20. For the final experiments, we decided to use deterministic decoding which is non-parametric and conceptually more suitable for D2T generation.
Prefixing the output makes parsing easier.
Even with variations of a “generate only the output” instruction appended to the prompt, the models (especially Llama2) tended to first confirm the request. For that reason, we decided to prefix the input for all the models with “Sure! Here is the {output_type}: "”. The opening quote at the end of the prefix allowed us to robustly parse the text simply by stripping the closing quote from the model output.
The outputs are fluent, but inaccurate.
We observed that the vast majority of model outputs were gramatically and stylistically correct, capturing the output type specified in the prompt. However, we also noticed that the outputs contained many factual errors (even after emphasizing the focus on factual accuracy in the prompt, see Figure 2). This observation lead us to evaluate the model outputs using token-level annotations focused on semantic accuracy errors Reiter and Thomson (2020).
3.4 Final Experiments
Taking the observations in §3.3 into account, we proceeded to generate the outputs on the test set of Quintd-1 for token-level error analysis. We first preprocessed the data as mentioned: we stripped out unnecessary fields, added units, and downsampled the data to fit the context. For all the models mentioned in §3.2, we used the prompt in Figure 2 and deterministic decoding with a maximum length of 512 tokens.
For comparsion, we also generated outputs for the same inputs and identical prompts with GPT-3.5 (gpt-3.5-turbo-1106). [^4] We did not include GPT-3.5 in our preliminary experiments since closed models are not our focus. We also did not use GPT-4 because we reserve the model for evaluation, see §4.1. Given that the decoding parameters provided by the OpenAI API are fairly limited, we only tried to bring the setup closer to deterministic decoding by fixing the temperature and seed to 00.
4 Evaluation
For evaluation, we focus on semantic accuracy errors. We compare the generated texts to the input data, looking for parts of texts which are not faithful to the input data. We annotate the errors on the token level, considering all the tokens in the output text as potential sources of errors.
Regarding the error taxonomy, we settled on four error categories: INCORRECT, NOT_CHECKABLE, MISLEADING, and OTHER. The taxonomy is inspired by the methodology discussed in Thomson and Reiter (2020) and Thomson et al. (2023). To keep the annotation tractable, we decided not to distinguish between fine-grained categories (e.g. incorrect name vs. incorrect number). The descriptions of our error categories, as presented in the instructions for annotation, are included in Table 2.
Error Description
INCORRECT
The fact in the text contradicts the data.
NOT_CHECKABLE
The fact in the text cannot be checked given the data.
MISLEADING
The fact in the text is misleading in the given context.
OTHER
The text is problematic for another reason, e.g. grammatically or stylistically incorrect, irrelevant, or repetitive.
Example
data
Nokia 3310 | color: black, blue, grey | display: 320x240px
text
Nokia 3310 is produced in Finland and features a 320x320 display. It is available in black color. The data seem to provide only partial information about the phone.
Table 2: Categories of errors annotated in our evaluation and an example demonstrating the error types. See Appendix C for explanation about individual errors in the example.
We employ two complementary evaluation schemes:
  • $\mathcal{E}_{\text{gpt}}$: an automatic metric based on GPT-4 (§4.1),
  • $\mathcal{E}_{\text{hum}}$: human evaluation based on crowdsourcing (§4.2).
These two schemes are based on similar instructions and produce (nearly) equivalent outputs. The main idea in introducing multiple schemes is to compensate for shortcomings of each approach and thus increase the replicability and robustness of our results.
4.1 GPT-4-based Evaluation
LLM-based metrics can be customized for a particular task without the need for training data. For our experiments, we employ a metric based on GPT-4 (gpt-4-1106-preview, OpenAI, 2023a), which was shown to be superior in following fine-grained instructions compared to other LLMs, and to be having high correlations with human judgement on evaluating generated texts Zhao et al. (2023); Sottana et al. (2023); Kocmi and Federmann (2023a, b).
$\mathcal{E}_{\text{gpt}}$ is instantiated using a prompt and a system message describing the task. We instruct the model to produce a JSON output with sequentially ordered errors using the following format:
  "errors": [{
    "reason": [REASON],
    "text": [TEXT_SPAN],
    "type": [ERROR_CATEGORY]
Note that we require that the model first generates the reason for the error. Besides that we get a free-form text reasoning for each error without excessive extra costs, our cursory observations suggest that this leads to more precise outputs.
Concerning the alignment of the model outputs with the original text, we perform matching on TEXT_SPAN. We ensure that the model response is a valid JSON using OpenAI’s response_format parameter. See Appendix B for more details about the metric, including the prompt and the system message.
4.2 Human-based Evaluation
An automatic metric based on a closed LLM makes the evaluation potentially non-reproducibile and biased Kocmi and Federmann (2023a); Wang et al. (2023b), for which we compensate by obtaining annotations from human annotators.
For the human annotation metric $\mathcal{E}_{\text{hum}}$, we prepared a custom web interface based on TabGenie Kasner et al. (2023a), where an annotator is able to annotate a text span with a selected error category. We created custom visualizations for each data format. Unlike with $\mathcal{E}_{\text{gpt}}$, we did not ask the crowdworkers for free-form reasoning about the errors since that would make the annotation more complex.
We hired annotators on the Prolific crowdsourcing platform. [^5] https://prolific.com In total, we hired 100 annotators, each annotating 20 examples (4 model outputs for each of the 5 domains). We selected annotators with at least 10 completed tasks and 100% approval rate, having English as their primary language. We paid the annotators £9 per hour, according to the platform’s recommendations. The median time for completing the annotations was 47 minutes. See Appendix C for the instructions for the annotators and the annotation interface and Appendix E for the data visualizations.
5 Results and Discussion
A summary of the token-level annotations is in Table 5 and 5, with detailed results per domain provided in Appendix F.
Table 3: The average numbers of errors per output (lower is better) based on GPT-4 ($\mathcal{E}_{\text{gpt}}$) and human annotators ($\mathcal{E}_{\text{hum}}$). We also include the average number of tokens per output in the rightmost column. The results of the best open LLM are emphasized.
Figure 5: The instructions given to the human annotators.
Incorrect Not Checkable Misleading Other All categories
$\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ Tok.
Llama2 2.46 1.50 0.90 1.21 0.20 0.23 0.13 0.10 3.70 3.05 83.8
Mistral 2.80 1.95 0.52 1.12 0.37 0.44 0.11 0.25 3.80 3.76 114.9
Zephyr 2.50 1.41 0.40 0.79 0.39 0.20 0.06 0.16 3.35 2.56 98.0
[…]
Incorrect Not Checkable Misleading Other All categories
$\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$
Llama2 0.78 0.52 0.46 0.57 0.15 0.17 0.09 0.08 0.92 0.85
Mistral 0.78 0.53 0.32 0.49 0.23 0.20 0.08 0.13 0.93 0.80
Zephyr 0.77 0.46 0.26 0.42 0.27 0.16 0.04 0.11 0.88 0.74
[…]
Table 4: The ratio of outputs containing at least one error (lower is better) based on GPT-4 ($\mathcal{E}_{\text{gpt}}$) and human annotators ($\mathcal{E}_{\text{hum}}$). The results of the best open LLM are emphasized.
5.1 How Accurate Are the Model Outputs?
Depending on the model, between 74-85% of examples contain an error according to $\mathcal{E}_{\text{hum}}$, suggesting that open LLMs make semantic errors very often. According to $\mathcal{E}_{\text{gpt}}$, the number is as high as 88-93%.
The most common error type is INCORRECT. As shown in Table 5, all the open LLMs make more than two statements contradicting the data per output on average. The NOT_CHECKABLE errors are also relatively common: more than one per output on average according to $\mathcal{E}_{\text{hum}}$, and at least one being present in more than 26% of examples according to both metrics.
The results vary widely according to the domain (see Appendix F). For example, the outputs in wikidata contain much more NOT_CHECKABLE errors on average (1.54 per output according to $\mathcal{E}_{\text{hum}}$) than INCORRECT errors (0.12 per output according to $\mathcal{E}_{\text{hum}}$), suggesting that with simpler inputs, the models tend towards extrinsic hallucinations. The openweather domain seems to be the most complex with the longest outputs (~164 tokens), more than 8 errors in the output on average and >90% of outputs containing an error.
The differences between the open LLMs are not major. Out of the open LLMs, Zephyr has the best results across categories and metrics, followed by Llama2. However, the outputs of Mistral are longer on average, leaving more space for errors. GPT-3.5 (which we consider separately) does generally better according to both $\mathcal{E}_{\text{gpt}}$ and $\mathcal{E}_{\text{hum}}$, although still makes an error in 60-75% of examples (2 errors per example on average). In general, the results show that LLMs make too many semantic errors to be usable in practice for D2T generation in zero-shot setting.
5.2 Do Evaluation Methods Agree?
To quantify the agreement of our evaluation metrics, we computed the Pearson correlation coefficient between the error counts on the level of tokens, examples, and domains (see Appendix D for details). The correlation on the level of domains ($r_{\text{domain}}=0.92$) shows that $\mathcal{E}_{\text{gpt}}$ and $\mathcal{E}_{\text{hum}}$ correlate well for high-level ranking of the models. However, the correlation is weaker on the level of individual examples ($r_{\text{example}}=0.52$), and even weaker on the level of tokens ($r_{\text{token}}=0.25$).
In Table 5, we show the percentage of tokens marked by individual metrics. While both metrics mark around 21% of tokens as errorneous, they agree on the specific tokens in less than 6%, suggesting that the error spans may be subject to interpretation. We conclude while the details regarding error spans and categories may vary, the metrics generally agree on accuracy of model outputs for a given set of examples.
$\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ + $\mathcal{E}_{\text{hum}}$
Incorrect 0.135 0.099 0.040
Not checkable 0.039 0.076 0.016
Misleading 0.022 0.022 0.001
Other 0.008 0.018 0.001
All categories 0.204 0.214 0.059
Table 5: The ratio of tokens marked as errorneous by GPT-4 ($\mathcal{E}_{\text{gpt}}$), human annotators ($\mathcal{E}_{\text{hum}}$), and both metrics at the same time ($\mathcal{E}_{\text{gpt}}$ + $\mathcal{E}_{\text{hum}}$).
5.3 Recommendations and Directions
Forget fluency, solve accuracy.
The output of LLMs is mostly satisfactory regarding the style, format, and purpose of the text. However, the amount of semantic errors remains very high. Improving semantic accuracy of the models Li et al. (2022); Schmidtová (2023), along with new model-based evaluation metrics Liu et al. (2023); Xu et al. (2023), could thus help to bring improve LLM-based D2T generation systems where it is most needed.
Use efficient models.
The memory issues with long context, making few-shot experiments infeasible, can potentially be solved by using more efficient long-context models equipped with Flash Attention Dao et al. (2022) and fast inference libraries such as exllama [^6] https://github.com/turboderp/exllamav2 or llama.cpp. [^7] https://github.com/ggerganov/llama.cpp A potentially suitable open LLM in this respect is the long-context Llama2 (Xiong et al., 2023; model to be released).
Test the models in the wild.
Except for using an ad-hoc dataset of real-world data as we did in our work, ecological validity of D2T evaluation can be also ensured by continous evaluation with human users Zheng et al. (2023) and evaluating real-world impact of the systems Reiter (2023).
Multilinguality is an opportunity.
With the recent efforts in extending D2T generation to low-resource languages Cripwell et al. (2023), multilingual D2T generation with open LLMs seems a promising direction. Although we did not go beyond English, initial steps were already done by works such as Lorandi and Belz (2023).
Be careful about subtle bugs.
During our preliminary experiments, we uncovered subtle bugs in API calls such as incorrect instruction templates [^8] https://huggingface.co/docs/transformers/chat_templating or involuntary input truncation. With the apparent ease of API access and robustness of LLMs, such bugs could go unnoticed and artificially skew the model performance.
6 Related Work
6.1 D2T Generation Tasks
Weather Forecasts
First attempts for generating weather forecasts include template-based and statistical approaches Belz (2005, 2008); Angeli et al. (2010) for the Sumtime-meteo and WeatherGov datasets Sripada et al. (2002); Liang et al. (2009). More recently, Balakrishnan et al. (2019) introduced a weather forecast dataset with tree-structured meaning representations. Our weather forecasts are less structured and based on a 5-day weather outlook.
Product Descriptions
Our phone specifications are closest to Wen et al. (2015, 2016) who introduced a dataset for generating descriptions of laptops and TVs. Their solution was based on recurrent neural networks, although templates remained a go-to approach for the task Wang et al. (2017). Recently, Shao et al. (2021) and Koto et al. (2022) also proposed specialized architectures based on pretrained language models for the data from big e-commerce platforms.
Sport Reports
All the D2T generation datasets from the Rotowire family Wiseman et al. (2017); Wang (2019), including SportSett:Basketball Thomson et al. (2021), and ESPN-NBA Nie et al. (2018) focus on generating basketball reports. Along with MLB Puduppully et al. (2019b), these datasets belong among the most challenging D2T datasets, attracting various neural-based solutions Puduppully et al. (2019a, 2022); Puduppully and Lapata (2021); Rebuffel et al. (2020). We use instead simpler data covering ice hockey game summaries.
Chart Captions
Following the early rule-based approaches Demir et al. (2008, 2012), the approaches for chart captioning recently tackle large-scale datasets from data analytic institutions Obeid and Hoque (2020); Kantharaj et al. (2022). We focus on one of the tasks from Sharma et al. (2021), which is generating descriptions of time series in the health domain.
Entity Descriptions
The task of generating descriptions for a knowledge graph has been covered extensively in D2T generation (Gardent et al., 2017; Ferreira et al., 2020; Agarwal et al., 2021; Chen et al., 2020; Ribeiro et al., 2020, inter alia). Our task is to describe an entity provided a list of its properties, which is closely related to generating entity descriptions from Wikipedia infotables Lebret et al. (2016).
6.2 D2T Generation with LLMs
Recent works have focused on exploring capabilities of closed LLMs on existing D2T generation datasets. Axelsson and Skantze (2023) evaluated GPT-3.5 OpenAI (2023b) on WebNLG, along with Yuan and Färber (2023), who also tested the model on the AGENDA dataset Koncel-Kedziorski et al. (2019). Both works found out that regardless of potential data contamination, the LLMs rank behind state-of-the-art finetuned models on automatic metrics. Zhao et al. (2023) tested closed models on modified table-to-text generation datasets and found out that in terms of faithfulness, GPT-4 can outperform state-of-the-art models.
6.3 Beyond Reference-Based Metrics
Many works have recently investigated the potential of using LLMs for automatic reference-free evaluation of generated text, generally achieving high correlations with human judgement Zhao et al. (2023); Sottana et al. (2023); Kocmi and Federmann (2023a, b); Chiang and Lee (2023); Wang et al. (2023a); Fu et al. (2023). However, they also voice concerns about its non-reproducibility Kocmi and Federmann (2023a) and potential bias of these models Wang et al. (2023b).
Holtzman et al. (2023) suggest that the research on LLMs should move away from reporting benchmark scores, investigating model behaviors instead. In this vein, Upadhyay and Massie (2022) analysed the ability of models of producing different types of content in D2T generation. Regarding human evaluation, Thomson and Reiter (2020) proposed a protocol for reference-free token-level annotation of complex D2T generation output.
7 Conclusion
We provided an exploratory study into D2T generation with open LLMs. We proposed new directions for D2T generation, including using ad-hoc test sets, data in common formats, and reference-free evaluation. By a combination of GPT-4-based metric and human evaluation, we evaluated the performance of LLMs on five domains, providing token-level annotations of model outputs across five domains, and recommendations for future directions in D2T generation.
Acknowledgements
This work was supported by the European Research Council (Grant agreement No. 101039303, NG-NLG) and Charles University project SVV 260 698. It used resources of the LINDAT/CLARIAH-CZ Research Infrastructure (Czech Ministry of Education, Youth, and Sports project No. LM2018101).
Limitations
In our work, we do not include comparison to other D2T generation approaches. The main reason is that our benchmark is reference-free, while large majority of prior approaches are based on models finetuned on reference outputs. However, we believe that our work still satisties our main goal of providing insights into behaviors of open LLM models on D2T generation.
We acknowledge that reference-free metrics currently have various shortcomings, including reliance on closed models or specific human annotation protocols, leading to limited replicability, and high price of execution. None of the approaches also produces flawless outcomes and have only moderate correlations with each other. We believe that these shortcomings will be adressed in the future with open model-based metrics.
Our choice of models is limited to 7B-parameter open LLMs due to our limited computational resources. Also, unlike some other LLMs such as GPT-Neo Black et al. (2022) or BLOOM BigScience Workshop et al. (2022), the models we used do not disclose the data they were trained on. For this reason, we find it ever more important to test the models on benchmarks whose labels could have not been included in their training data.
The approaches based on LLMs may produce factually incorrect information. Any text produced by the LLMs therefore need to be carefully examined, and no decisions should be based on the generated text alone.
Appendix A Quintd Data
Here we describe the data sources we include in the Quintd collection tool and the procedure of collecting the Quintd-1 benchmark. To replicate the data collection, please refer to the scripts we provide.
A.1 Selection of Data Sources
When selecting the data sources, we had the following desiderata:
  • Data needs to be publicly available.
  • Data needs to represent a common data-to-text task.
  • Data needs to be in a common format (or straightforwardly transformable to one).
We settled on the data sources described in Appendix 6. All the sources can be accessed using an API. Note that some of the APIs have access limits, either for the requests made from a single account per day, or for a number of requests from an IP address within a time window. However, these limits do not severely limit the data collection process on the scale we use here.
A.2 Data Collection
Table 6 summarizes the sources of data and output types for each domain.
Domain Id Source Output type
openweather OpenWeather five-day weather forecast
gsmarena GSMArena product description
ice_hockey RapidAPI ice hockey game summary
owid OurWorldInData chart caption
wikidata Wikidata entity description
Table 6: The sources of data and output types for individual domains in Quintd.
A.2.1 OpenWeather
OpenWeather (OpenWeatherMap.org) is an online service which provides global weather data via web interface and API. The API responses are in the JSON format documented at the official website. For our experiments, we used the forecast5 API which allows to download a 5-day forecast with 3-hour resolution for any location specified by its GPS coordinates.
The free tier is limited to 1,000 API calls per day, which is enough to download our whole test set in one bulk. However, at the time of experiments, the free API only allowed to download the data for the time when the request was made. At the time of writing, OpenWeather is pushing a new One Call API 3.0 which allows to download weather data for any timestamp, but only 4 days ahead (instead of 5). These restrictions somehow limit the replicability of our Quintd-1 dataset (at least with the free API), but do not limit downloading a new batch of data with a similar format.
For the Quintd-1 dataset, we randomly sampled 100 cities for each split from the list of cities with a population over 1000 and used their coordinates in the queries to OpenWeather API. All the data forecasts were downloaded on Jan 3, 2024.
A.2.2 GSMArena
GSMArena is a website providing specifications and reviews for mobile devices. For downloading the data, we used the unofficial gsmarena-api tool which returns the data in a JSON format. Note that GSMArena imposes limitations on the number of requests per IP address, which may induce delays when downloading a larger amount of data.
To create a balanced sample, we downloaded detailed specification of 10 products from each available brand, and randomly selected 100 products for each split from the downloaded set.
A.2.3 RapidAPI Ice Hockey
RapidAPI is a service which provides API access to data from multiple domains, including sport, finance, entertainment, and others. Most of the APIs are provided in a freemium mode, i.e. with a limited number of API calls per day.
For Quintd, we selected the IceHockeyAPI (popularity 9.1 / 10) which provides access to ice hockey games from world top leagues. Our choice was influenced by our own personal preferences, combined with the desire to cover a sport which has not been covered previously in sports report generation.
We used the matches endpoint which returns high-level details about a game. Note that the API allows only 50 requests per day, but that does not limit the data collection since the endpoint returns all the games played on a particular day in a single request. We downloaded the games played on 27 November 2023 for the development set (184 games) and 29 November 2023 for the test set (216 games), taking a random sample of 100 for each split.
A.2.4 OurWorldInData
OurWorldInData is a public database and web interface for data about world developments in various domains and sources. We used the official API (currently experimental) which is accessible through the Python package owid-catalog. The package allows to access individual CSV tables as Pandas dataframes.
For our data collection, we decided to limit ourselves to time series, i.e. a single column with values changing over time. Besides the simplicity of visualizing such a chart (which is used by human annotators for checking the correctness of the output), there is also a clear goal for the target chart description: describing the developments of a value over time. We also limited ourselves to the health domain. In particular, we selected the tables COVID data (columns new_cases_smoothed_per_million, new_tests_smoothed_per_thousand, people_vaccinated_per_hundred, reproduction_rate, and positive_rate) and Life expectancy data (column life_expectancy_0).
We downloaded the data for all countries with non-empty entries in the table, taking a random sample of 100 examples for each split. On model input, we formatted the data for each time series as two-column CSV, including the title, the description, and the unit for each example as a comment (#) at the beginning of the input.
A.2.5 Wikidata
Wikidata is a large open-source knowledge graph, containing factual information about entities and their properties. Wikidata provides access through an official API, but we instead decided to extract our data using the wikidatasets Boschin and Bonald (2019) Python library, which provides access to preprocessed properties of entities from particular domains. It allowed us to avoid crawling and filtering the knowledge graph and its offline processing made the data collection faster. [^9] All the entities and properties are linked with an identifier to the Wikidata database, making the process also replicable through the official API.
For our dataset, we selected the entities from the companies, countries, films, and humans domains. For each entity, we randomly extracted between 2 to 10 properties in the knowledge graph. We extracted up to 100 subgraphs for each domain and took a random sample of 100 subgraphs for each split. On model input, we formatted each subgraph as a simple Markdown-formatted text snippet, using the entity as a title and including a bullet point for each key-value pair.
Appendix B GPT-4 Evaluation
We used the prompt in Figure 4 for instantiating the GPT-4-based metric. [^10] Note that the example in the prompt differs from the example used for human annotators (see Figure 5). We revised the example to be more instructive but we were not able to re-run the GPT-4 evaluation due to our limited budget. We ensured that the output is a valid JSON using a recently introduced parameter response_format in the OpenAI API. At the price of \$0.01 per 1k input tokens and \$0.03 per 1k generated tokens, the evaluation process cost approximately \$45 in total.
B.1 Aligning the Errors
For aligning the errors with the original text, we perform string matching on the text span decoded by GPT-4 in the TEXT_SPAN field. In our preliminary experiments, this method proved to be more robust than either asking for start and end indices of the error span (which would rely on the model’s ability to count characters) or performing sequence tagging on the copy of the input (which would rely on the model’s ability to perfectly copy the input).
We tried to respect monotonic ordering of text spans but fell back to full-text search if the span is not found following the previous one. We consider this approach successful since matching completely failed only in a minority of cases (137 out of 6927). Based on our manual examination, these mostly include cases where GPT-4 tried to suggest a missing piece of text as an error, or did not manage to copy the input text verbatim.
Appendix C Human Evaluation
As described in §4.2, we set up the human evaluation campaign on Prolific. To make the data more accessible to the annotators, we created custom data visualizations for each domain. For the data in openweather and owid, we used interactive graphs from Highcharts.com, and we manually created the tables for other domains. You can find the full instructions for human annotators in Figure 5 and the examples of data visualizations in Appendix E.
Appendix D Metric Correlation
The Pearson correlation coefficients (§5.2) were computed using two lists (for $\mathcal{E}_{\text{hum}}$ and $\mathcal{E}_{\text{gpt}}$) as follows (note that each error category was considered separately):
  • For $r_{\text{domain}}$, we concatenated the average error counts per domain (see Table 5).
  • For $r_{\text{example}}$, we concatenated the count of errors per example.
  • For $r_{\text{token}}$, we concatenated the binary indicators marking an error per token.
Appendix E Examples
Here we present an example of inputs and model outputs (along with annotations) for each domain:
  • openweather: Figure 7 (in) and Table 7 (out),
  • gsmarena: Figure 8 (in) and Table 7 (out),
  • ice_hockey: Figure 9 (in) and Table 9 (out),
  • owid: Figure 10 (in) and Table 10 (out),
  • wikidata: Figure 11 (in) and Table 11 (out).
Note that the graphs for openweather and owid are interactive when accessed through the web interface.
Appendix F Full Results
Here we include the tables with results for individual domains:
  • Table 5 presents the average numbers of errors per output separately for each domain (the aggregated results are in Table 5),
  • Table 5 presents the ratio of outputs containing at least one error separately for each domain (the aggregated results are in Table 5).
System Message
Prompt
Figure 4: The prompt we used for the GPT-4 evaluation metric.
By continuing, you agree that the anonymous annotations will be made available to the research community to help evaluating and improving data-to-text generation systems based on large language models.
Figure 6: The annotation interface for human annotators.
Figure 7: An example of an input from openweather (see the outputs in Table 7).
Model
Human annotations ($\mathcal{E}_{\text{hum}}$)
GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
Llama2
The weather in Niemegk, Germany, will be cloudy and rainy for the next few days. On January 3rd, the temperature will be around 9.55°C, with a feels-like temperature of 6.59°C. The weather will be light rain, with a wind speed of 6.24 m/s and a gust of 12.1 m/s. The rain will last for 3 hours, and the pressure will be 991 hPa.
The weather in Niemegk, Germany, will be cloudy and rainy for the next few days. On January 3rd, the temperature will be around 9.55°C, with a feels-like temperature of 6.59°C. The weather will be light rain, with a wind speed of 6.24 m/s and a gust of 12.1 m/s. The rain will last for 3 hours, and the pressure will be 991 hPa.
Mistral
From January 3rd to January 7th, the city of Niemegk in Germany will experience a range of weather conditions. On January 3rd, the temperature will be 9.55°C with a feels like temperature of 6.59°C. There will be light rain with a 3-hour precipitation of 1.06mm. The wind speed will be 6.24m/s with a gust of 12.1m/s. On January 4th, the temperature will be 8.2°C with a feels like temperature of 5.64°C. There will be light rain with a 3-hour precipitation of 1.94mm. The wind speed will be 4.31m/s with a gust of 11.6m/s. On January 5th, the temperature will be 6.37°C with a feels like temperature of 2.25°C. There will be light rain with a 3-hour precipitation of 1.13mm. The wind speed will be 7.07m/s with a gust of 15.51m/s. On January 6th, the temperature will be 2.17°C with a feels like temperature of -3.55°C. There will be overcast clouds with a 3-hour precipitation of 0.31mm. The wind speed will be 3.54m/s with a gust of 7.11m/s. On January 7th, the temperature will be -0.89°C with a feels like temperature of -0.89°C. There will be overcast clouds with a 3-hour precipitation of 0.31mm. The wind speed will be 0.63m/s with a gust of 0.7m/s.
From January 3rd to January 7th, the city of Niemegk in Germany will experience a range of weather conditions. On January 3rd, the temperature will be 9.55°C with a feels like temperature of 6.59°C. There will be light rain with a 3-hour precipitation of 1.06mm. The wind speed will be 6.24m/s with a gust of 12.1m/s. On January 4th, the temperature will be 8.2°C with a feels like temperature of 5.64°C. There will be light rain with a 3-hour precipitation of 1.94mm. The wind speed will be 4.31m/s with a gust of 11.6m/s. On January 5th, the temperature will be 6.37°C with a feels like temperature of 2.25°C. There will be light rain with a 3-hour precipitation of 1.13mm. The wind speed will be 7.07m/s with a gust of 15.51m/s. On January 6th, the temperature will be 2.17°C with a feels like temperature of -3.55°C. There will be overcast clouds with a 3-hour precipitation of 0.31mm. The wind speed will be 3.54m/s with a gust of 7.11m/s. On January 7th, the temperature will be -0.89°C with a feels like temperature of -0.89°C. There will be overcast clouds with a 3-hour precipitation of 0.31mm. The wind speed will be 0.63m/s with a gust of 0.7m/s.
Zephyr
Over the next five days, the weather in Niemegk, Germany will be characterized by fluctuating temperatures and varying precipitation. On Thursday, January 3, light rain and a low temperature of 9.55°C can be expected, with a chance of rain continuing into Friday. Snowfall is possible on Friday, with a low temperature of -0.89°C. Saturday will be overcast with a low temperature of 0.03°C, and light rain is forecasted for Saturday evening. Sunday will be overcast as well, with a low temperature of -2.23°C. Monday will see a drop in temperatures, with a low of -4.07°C and scattered clouds. Temperatures will continue to fall throughout the week, with lows of -5.19°C and -9.51°C on Tuesday and Wednesday, respectively. Wind speeds will vary from 3.1 to 15.51 meters per second, with gusts reaching up to 15.51 meters per second on Monday.
Over the next five days, the weather in Niemegk, Germany will be characterized by fluctuating temperatures and varying precipitation. On Thursday, January 3, light rain and a low temperature of 9.55°C can be expected, with a chance of rain continuing into Friday. Snowfall is possible on Friday, with a low temperature of -0.89°C. Saturday will be overcast with a low temperature of 0.03°C, and light rain is forecasted for Saturday evening. Sunday will be overcast as well, with a low temperature of -2.23°C. Monday will see a drop in temperatures, with a low of -4.07°C and scattered clouds. Temperatures will continue to fall throughout the week, with lows of -5.19°C and -9.51°C on Tuesday and Wednesday, respectively. Wind speeds will vary from 3.1 to 15.51 meters per second, with gusts reaching up to 15.51 meters per second on Monday.
GPT-3.5
Over the next five days, the weather in Niemegk, Germany is expected to be quite varied. On January 4th, there will be light rain with temperatures ranging from 2.17°C to 9.55°C. The following day, January 5th, there will be light snow with temperatures dropping to as low as -0.89°C. On January 6th, the weather will be overcast with temperatures ranging from -0.32°C to 1.69°C. The trend continues on January 7th with overcast clouds and temperatures dropping to as low as -2.51°C. Finally, on January 8th, there will be scattered clouds with temperatures ranging from -5.19°C to -3.67°C. Throughout this period, wind speeds will vary between 0.63 m/s and 8.12 m/s, and the pressure will range from 991 hPa to 1034 hPa.
Over the next five days, the weather in Niemegk, Germany is expected to be quite varied. On January 4th, there will be light rain with temperatures ranging from 2.17°C to 9.55°C. The following day, January 5th, there will be light snow with temperatures dropping to as low as -0.89°C. On January 6th, the weather will be overcast with temperatures ranging from -0.32°C to 1.69°C. The trend continues on January 7th with overcast clouds and temperatures dropping to as low as -2.51°C. Finally, on January 8th, there will be scattered clouds with temperatures ranging from -5.19°C to -3.67°C. Throughout this period, wind speeds will vary between 0.63 m/s and 8.12 m/s, and the pressure will range from 991 hPa to 1034 hPa.
Table 7: An example of the model outputs on openweather (see the input in Figure 7).
Figure 8: An example of an input from gsmarena (see the outputs in Table 8).
Model
Human annotations ($\mathcal{E}_{\text{hum}}$)
GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
Llama2
The Gigabyte GSmart Mika M3 is a smartphone that was released in 2014. It has a 5-inch IPS LCD display with a resolution of 720x1280 pixels and a 13MP main camera. The device also has an 8MP front camera and is powered by a Mediatek MT6582 chipset with 1GB of RAM and 8GB of internal storage. It has dual SIM support and is available in Navy Blue.
The Gigabyte GSmart Mika M3 is a smartphone that was released in 2014. It has a 5-inch IPS LCD display with a resolution of 720x1280 pixels and a 13MP main camera. The device also has an 8MP front camera and is powered by a Mediatek MT6582 chipset with 1GB of RAM and 8GB of internal storage. It has dual SIM support and is available in Navy Blue.
Mistral
The GSmart Mika M3 is a smartphone developed by Gigabyte. It was released in November 2014 and is powered by the Android 4.4.2 operating system. The device features a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels, which translates to 294 pixels per inch. The smartphone is equipped with a 13-megapixel rear camera with autofocus and LED flash, as well as an 8-megapixel front-facing camera. It is powered by a Mediatek MT6582 chipset with a quad-core 1.3 GHz Cortex-A7 CPU and a Mali-400MP2 GPU. The device has 8GB of internal storage and 1GB of RAM, with the option to expand storage via a microSDHC card. The GSmart Mika M3 also features Wi-Fi, Bluetooth, GPS, and NFC connectivity, as well as a 3.5mm headphone jack and a microUSB 2.0 port for charging and data transfer. The device is available in Navy Blue and is powered by a 1900mAh Li-Ion battery, which provides up to 150 hours of stand-by time and up to 4 hours and 30 minutes of talk time.
The GSmart Mika M3 is a smartphone developed by Gigabyte. It was released in November 2014 and is powered by the Android 4.4.2 operating system. The device features a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels, which translates to 294 pixels per inch. The smartphone is equipped with a 13-megapixel rear camera with autofocus and LED flash, as well as an 8-megapixel front-facing camera. It is powered by a Mediatek MT6582 chipset with a quad-core 1.3 GHz Cortex-A7 CPU and a Mali-400MP2 GPU. The device has 8GB of internal storage and 1GB of RAM, with the option to expand storage via a microSDHC card. The GSmart Mika M3 also features Wi-Fi, Bluetooth, GPS, and NFC connectivity, as well as a 3.5mm headphone jack and a microUSB 2.0 port for charging and data transfer. The device is available in Navy Blue and is powered by a 1900mAh Li-Ion battery, which provides up to 150 hours of stand-by time and up to 4 hours and 30 minutes of talk time.
Zephyr
The Gigabyte GSmart Mika M3 is a smartphone that boasts a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels. It runs on Android 4.4.2 (KitKat) and is powered by a quad-core 1.3 GHz Cortex-A7 processor with a Mali-400MP2 GPU. The device comes with 8GB of internal storage and 1GB of RAM, which can be expanded via a dedicated microSDHC card slot. The primary camera is a 13-megapixel autofocus lens with an LED flash, while the secondary camera is an 8-megapixel lens. Connectivity options include Wi-Fi, Bluetooth, GPS, FM radio, and microUSB 2.0. The smartphone is equipped with a 1900mAh removable Li-Ion battery that provides up to 4 hours of talk time and up to 150 hours of standby time. The device is available in Navy Blue and supports dual SIM cards (Mini-SIM/ Micro-SIM).
The Gigabyte GSmart Mika M3 is a smartphone that boasts a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels. It runs on Android 4.4.2 (KitKat) and is powered by a quad-core 1.3 GHz Cortex-A7 processor with a Mali-400MP2 GPU. The device comes with 8GB of internal storage and 1GB of RAM, which can be expanded via a dedicated microSDHC card slot. The primary camera is a 13-megapixel autofocus lens with an LED flash, while the secondary camera is an 8-megapixel lens. Connectivity options include Wi-Fi, Bluetooth, GPS, FM radio, and microUSB 2.0. The smartphone is equipped with a 1900mAh removable Li-Ion battery that provides up to 4 hours of talk time and up to 150 hours of standby time. The device is available in Navy Blue and supports dual SIM cards (Mini-SIM/ Micro-SIM).
GPT-3.5
The GSmart Mika M3, also known as the Gigabyte GSmart Mika M3, is a smartphone that was announced in October 2014 and released in November of the same year. It features a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels and is powered by a Quad-core 1.3 GHz Cortex-A7 processor. The phone comes with 1GB of RAM and 8GB of internal storage, which can be expanded via a microSDHC card slot. The main camera is 13MP with LED flash, while the selfie camera is 8MP. The phone runs on Android 4.4.2 (KitKat) and is equipped with a 1900mAh removable battery. Other features include dual SIM capability, Wi-Fi, Bluetooth, GPS, FM radio, and microUSB 2.0. The GSmart Mika M3 is available in the color Navy Blue and has a sleek and compact design, measuring 144 x 70.5 x 8.3 mm and weighing 146g.
The GSmart Mika M3, also known as the Gigabyte GSmart Mika M3, is a smartphone that was announced in October 2014 and released in November of the same year. It features a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels and is powered by a Quad-core 1.3 GHz Cortex-A7 processor. The phone comes with 1GB of RAM and 8GB of internal storage, which can be expanded via a microSDHC card slot. The main camera is 13MP with LED flash, while the selfie camera is 8MP. The phone runs on Android 4.4.2 (KitKat) and is equipped with a 1900mAh removable battery. Other features include dual SIM capability, Wi-Fi, Bluetooth, GPS, FM radio, and microUSB 2.0. The GSmart Mika M3 is available in the color Navy Blue and has a sleek and compact design, measuring 144 x 70.5 x 8.3 mm and weighing 146g.
Table 8: An example of the model outputs on gsmarena (see the input in Figure 8).
Figure 9: An example of an input from ice_hockey (see the outputs in Table 9).
Model
Human annotations ($\mathcal{E}_{\text{hum}}$)
GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
Llama2
In the VHL 23/24 season, Kristall Saratov hosted HC Norilsk at their home arena. Kristall Saratov won the game with a final score of 1-5. Kristall Saratov’s home team scored one goal in the first period, while HC Norilsk scored two goals in the second and third periods. The game lasted for 60 minutes, with Kristall Saratov’s home team leading the game for the entire duration.
In the VHL 23/24 season, Kristall Saratov hosted HC Norilsk at their home arena. Kristall Saratov won the game with a final score of 1-5. Kristall Saratov’s home team scored one goal in the first period, while HC Norilsk scored two goals in the second and third periods. The game lasted for 60 minutes, with Kristall Saratov’s home team leading the game for the entire duration.
Mistral
On November 28th, 2023, the VHL 23/24 season game between Kristall Saratov and HC Norilsk took place. The game was played at the home arena of Kristall Saratov, with the home team winning 1-0. The game was played in three periods, with the first period ending in a 0-0 tie, the second period ending in a 2-3 score in favor of HC Norilsk, and the third period ending in a 1-0 score in favor of Kristall Saratov. The game ended in normal time, with a total score of 3-1 in favor of HC Norilsk.
On November 28th, 2023, the VHL 23/24 season game between Kristall Saratov and HC Norilsk took place. The game was played at the home arena of Kristall Saratov, with the home team winning 1-0. The game was played in three periods, with the first period ending in a 0-0 tie, the second period ending in a 2-3 score in favor of HC Norilsk, and the third period ending in a 1-0 score in favor of Kristall Saratov. The game ended in normal time, with a total score of 3-1 in favor of HC Norilsk.
Zephyr
In the VHL 23/24 season, Kristall Saratov hosted HC Norilsk. The first period saw two quick goals from Norilsk, with one in the second and third periods as well, resulting in a dominant 5-0 victory. Kristall Saratov managed to score a single goal in the third period, but it was not enough to overcome Norilsk’s lead. The final score was 1-5 in favor of HC Norilsk.
In the VHL 23/24 season, Kristall Saratov hosted HC Norilsk. The first period saw two quick goals from Norilsk, with one in the second and third periods as well, resulting in a dominant 5-0 victory. Kristall Saratov managed to score a single goal in the third period, but it was not enough to overcome Norilsk’s lead. The final score was 1-5 in favor of HC Norilsk.
GPT-3.5
In a match from the VHL 23/24 season, Kristall Saratov faced off against HC Norilsk. The game took place on November 28, 2023, at 15:30. Despite a goal from Kristall Saratov in the third period, HC Norilsk emerged victorious with a final score of 5-1. The away team dominated the game, scoring 2 goals in the first period and 3 goals in the second period. Kristall Saratov struggled to keep up, resulting in a clear win for HC Norilsk.
In a match from the VHL 23/24 season, Kristall Saratov faced off against HC Norilsk. The game took place on November 28, 2023, at 15:30. Despite a goal from Kristall Saratov in the third period, HC Norilsk emerged victorious with a final score of 5-1. The away team dominated the game, scoring 2 goals in the first period and 3 goals in the second period. Kristall Saratov struggled to keep up, resulting in a clear win for HC Norilsk.
Table 9: An example of the model outputs on ice_hockey (see the input in Figure 9).
Figure 10: An example of an input from owid (see the outputs in Table 10).
Model
Human annotations ($\mathcal{E}_{\text{hum}}$)
GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
Llama2
Life expectancy at birth in Djibouti has increased significantly over the past few decades, with a notable increase from 40.5665 years in 1950 to 62.3049 years in 2021.
Life expectancy at birth in Djibouti has increased significantly over the past few decades, with a notable increase from 40.5665 years in 1950 to 62.3049 years in 2021.
Mistral
This chart displays the life expectancy at birth in Djibouti from 1950 to 2021. The data shows a steady increase in life expectancy over the years, with a significant jump in the 1970s and 1980s. In 2021, the average life expectancy at birth in Djibouti was 62.3 years.
This chart displays the life expectancy at birth in Djibouti from 1950 to 2021. The data shows a steady increase in life expectancy over the years, with a significant jump in the 1970s and 1980s. In 2021, the average life expectancy at birth in Djibouti was 62.3 years.
Zephyr
This line graph illustrates the evolution of life expectancy at birth in Djibouti from 1950 to 2021. Initially, the average number of years that a newborn could expect to live was around 40.5 years in 1950, and it gradually increased over the following decades, reaching 62.3 years in 2021. The trend shows a steady improvement, with some fluctuations in the 1980s and 1990s. The data presented here is based on the sex- and age-specific death rates prevailing at the time of birth for a specific year in a given country, territory, or geographic area.
This line graph illustrates the evolution of life expectancy at birth in Djibouti from 1950 to 2021. Initially, the average number of years that a newborn could expect to live was around 40.5 years in 1950, and it gradually increased over the following decades, reaching 62.3 years in 2021. The trend shows a steady improvement, with some fluctuations in the 1980s and 1990s. The data presented here is based on the sex- and age-specific death rates prevailing at the time of birth for a specific year in a given country, territory, or geographic area.
GPT-3.5
The chart shows the life expectancy at birth in Djibouti from 1950 to 2021. The life expectancy has steadily increased over the years, from 40.57 years in 1950 to 62.30 years in 2021, with some fluctuations in between. The data reflects the average number of years a newborn could expect to live if exposed to the prevailing death rates at the time of birth.
The chart shows the life expectancy at birth in Djibouti from 1950 to 2021. The life expectancy has steadily increased over the years, from 40.57 years in 1950 to 62.30 years in 2021, with some fluctuations in between. The data reflects the average number of years a newborn could expect to live if exposed to the prevailing death rates at the time of birth.
Table 10: An example of the model outputs on owid (see the input in Figure 10).
Figure 11: An example of an input from wikidata (see the outputs in Table 11).
Model
Human annotations ($\mathcal{E}_{\text{hum}}$)
GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
Llama2
Driven Brands is an American franchise holding company that owns, operates, and franchises automotive services businesses in the United States. The company was founded in 1996 and is headquartered in Charlotte, North Carolina. It operates several brands, including Maaco, Meineke, Econo Lube, and CARSTAR.
Driven Brands is an American franchise holding company that owns, operates, and franchises automotive services businesses in the United States. The company was founded in 1996 and is headquartered in Charlotte, North Carolina. It operates several brands, including Maaco, Meineke, Econo Lube, and CARSTAR.
Mistral
Driven Brands is a business operating in the automotive services industry in the United States of America.
Driven Brands is a business operating in the automotive services industry in the United States of America.
Zephyr
Driven Brands is a business operating in the automotive services industry based in the United States of America. As an instance of the broader category of businesses, Driven Brands provides goods or services to customers in exchange for financial remuneration.
Driven Brands is a business operating in the automotive services industry based in the United States of America. As an instance of the broader category of businesses, Driven Brands provides goods or services to customers in exchange for financial remuneration.
GPT-3.5
Driven Brands is a business based in the United States of America, operating within the automotive services industry.
Driven Brands is a business based in the United States of America, operating within the automotive services industry.
Table 11: An example of the model outputs on wikidata (see the input in Figure 11).
Table 12: The average numbers of errors per output for each domain (lower is better). We also include the average number of tokens per output in the rightmost column. See Table 5 for aggregated results.
Incorrect Not Checkable Misleading Other All categories
$\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ Tok.
Llama2 openweather 5.77 4.02 1.94 2.06 0.34 0.52 0.43 0.16 8.48 6.76 166.7
gsmarena 1.88 0.65 0.37 0.89 0.33 0.19 0.11 0.09 2.69 1.82 90.3
ice_hockey 1.82 1.45 0.85 0.96 0.04 0.09 0.06 0.12 2.77 2.62 63.3
owid 2.18 1.25 0.22 0.58 0.22 0.28 0.01 0.07 2.63 2.18 60.3
wikidata 0.65 0.15 1.13 1.58 0.08 0.09 0.06 0.07 1.92 1.89 38.5
[…]
gsmarena 1.86 0.51 0.80 1.10 0.54 0.29 0.05 0.12 3.25 2.02 146.3
ice_hockey 1.91 1.44 0.72 0.89 0.13 0.14 0.05 0.13 2.81 2.60 92.4
owid 2.67 2.39 0.18 0.40 0.41 0.29 0.04 0.07 3.30 3.15 91.1
wikidata 0.40 0.10 0.32 0.71 0.27 0.13 0.32 0.38 1.31 1.32 51.0
[…]
gsmarena 1.72 0.34 0.93 1.04 0.85 0.23 0.03 0.10 3.53 1.71 142.8
ice_hockey 1.49 0.88 0.19 0.60 0.08 0.10 0.02 0.07 1.78 1.65 83.1
owid 2.51 1.63 0.20 0.52 0.43 0.22 0.01 0.10 3.15 2.47 85.2
wikidata 0.33 0.09 0.24 0.70 0.23 0.11 0.21 0.36 1.01 1.26 48.1
[…]
gsmarena 1.30 0.20 1.13 0.80 0.63 0.21 0.03 0.14 3.09 1.35 129.5
ice_hockey 0.72 0.81 0.12 0.48 0.05 0.07 0.00 0.08 0.89 1.44 84.4
owid 1.81 0.63 0.03 0.25 0.35 0.17 0.00 0.01 2.19 1.06 62.2
wikidata 0.14 0.05 0.13 0.35 0.11 0.06 0.06 0.05 0.44 0.51 35.7
Incorrect Not Checkable Misleading Other All categories
$\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$
Llama2 openweather 0.96 0.74 0.70 0.70 0.26 0.29 0.25 0.12 1.00 0.88
gsmarena 0.83 0.38 0.28 0.48 0.21 0.18 0.06 0.05 0.90 0.74
ice_hockey 0.77 0.69 0.49 0.50 0.04 0.07 0.05 0.10 0.93 0.88
owid 0.92 0.68 0.18 0.37 0.19 0.24 0.01 0.06 0.93 0.87
wikidata 0.42 0.12 0.63 0.79 0.07 0.09 0.06 0.05 0.83 0.87
[…]
gsmarena 0.89 0.33 0.45 0.53 0.33 0.18 0.05 0.11 0.99 0.72
ice_hockey 0.83 0.73 0.52 0.60 0.13 0.12 0.04 0.09 0.96 0.87
owid 0.91 0.74 0.17 0.27 0.29 0.24 0.04 0.06 0.98 0.85
wikidata 0.29 0.08 0.25 0.43 0.19 0.10 0.20 0.24 0.70 0.66
[…]
gsmarena 0.87 0.23 0.53 0.47 0.45 0.19 0.02 0.08 0.99 0.60
ice_hockey 0.79 0.54 0.16 0.38 0.08 0.10 0.02 0.06 0.84 0.75
owid 0.93 0.68 0.14 0.31 0.30 0.20 0.01 0.07 0.96 0.83
wikidata 0.28 0.06 0.20 0.48 0.21 0.07 0.15 0.22 0.60 0.64
[…]
gsmarena 0.69 0.15 0.68 0.42 0.34 0.13 0.03 0.11 0.93 0.56
ice_hockey 0.52 0.64 0.09 0.30 0.05 0.07 0.00 0.07 0.56 0.76
owid 0.90 0.42 0.02 0.13 0.30 0.14 0.00 0.01 0.92 0.57
wikidata 0.12 0.04 0.11 0.26 0.10 0.06 0.06 0.05 0.33 0.37
Table 13: The ratio of outputs containing at least one error for each domain (lower is better). See Table 5 for aggregated results.
Table 12: The average numbers of errors per output for each domain (lower is better). We also include the average number of tokens per output in the rightmost column. See Table 5 for aggregated results.
Table 3: The average numbers of errors per output (lower is better) based on GPT-4 ($\mathcal{E}_{\text{gpt}}$) and human annotators ($\mathcal{E}_{\text{hum}}$). We also include the average number of tokens per output in the rightmost column. The results of the best open LLM are emphasized.
Figure 5: The instructions given to the human annotators.